# Fractal geometry predicts dynamic differences in structural and functional connectomes

### v1

@@ 3.1 Structural connectomes
The graph-theoretical measures described in Section 2.2 were computed for all Structural connectomes. The mean and standard deviation for each of these measures are provided in Table 1. In Figure 5, we have the pairwise correlation between the graph-theoretical measures. Notice that, aside from $BC$ and $AS$, the other network measures show significant pairwise positive correlations with significance values $p<0.0007$.
Table 1: Network measures statistics for Structural connectomes: Betweenness Centrality (BC), Assortativity (AS), Clustering Coefficient (CC), Degree (D), Global Efficiency (GE), Local Efficiency (LE), Eigencentrality (EC), Transitivity (TR), Mean 3D Motifs 1 (M1) and 2 (M2). For the node-wise measures, the mean was first computed over all nodes. Then, mean $\mu$ and standard deviation $\sigma$ over the whole subject population are shown in the table for each of the measures used. The strengths of the correlations between these measures ($\rho$ values for Pearson correlations) are illustrated in Figure 5.
@@
The measures described in Section 2.4 were calculated for each Structural connectome. Their mean and standard deviation are reported in Table 2. To start with, Figure 12 readily conveys certain unifying and differentiating features across the Structural equi-M sets. First, we notice relatively broad differences in the size of the sets between subjects. These differences reflect relatively broad distributions in the values of the geometric landmarks and distances (for all diameters $d_{XY}$ that we calculated, the standard deviation is at least half of the mean). The one measure which has a significantly smaller relative variance is the ratio $\epsilon$. This is because the “eccentricity” of the sets remains relatively robust across the connectomes, with eccentricities tightly clustered around the average $\mu(\epsilon)=0.81$, regardless of size.
Table 2: Geometric measure statistics for Structural connectomes. Then mean $\mu$ and standard deviation $\sigma$ over the whole subject population were computed for each geometric measure. For the modes, the standard deviation $\sigma$ was computed by treating the real and imaginary parts as independent and calculating the standard deviation as the square root of the sum of the variances of the real and imaginary parts. The strengths of the correlations between these measures ($\rho$ values for Pearson correlations) are illustrated in Figure 5.
@@
Mode $m_{-2}$ $7.79$ $-2.59$ $6.33$
Mode $m_{-1}$ $-66.97$ $10.56$ $32.28$
Mode $m_{0}$ $-39.05$ $-0.09$ $20.15$
Mode $m_{1}$ $-10.52$ $-1.79$ $6.38$
Mode $m_{2}$ $-9.42$ $-3.15$ $6.19$
@@
As discussed in Section 2.3, we approach the collection of sets obtained from connectome data in a statistical way by of considering the distribution of equi-M sets over all connectomes under consideration. We did this by computing the equi-M set for each individual connectome and assigning to each point in the complex plane the fraction of individual equi-M sets that contain the point (effectively measuring how “likely” it is for the given point to be in any equi-M set generated by our data). Figure 4a illustrates this statistical equi-M set with the fraction represented by the color intensity, increasing from light to dark. To emphasize the underpinnings of this representation, we overlaid on it the contours of all equi-M sets for Structural connectomes (in gray), outlining the basis of how these statistics were computed. Consistent with our previous comment, all these contours exhibit a cusp on the right side along the real axis. Note, in particular, that the black center region is the $c$-locus that is common to the equi-M sets for all Structural connectomes.
Figure 4: Different approaches to the Structural connectome data using the equi-M set. (a) Statistical equi-M set (the light to dark gradient shows the fraction of equi-M sets that the respective point in $\mathbb{C}$ belongs to), together with the the collection of contours of all equi-M sets derived from Structural connectomes (overlaid in gray). (b) The equi-M set computed for the prototypical Structural connectome (i.e., the matrix computed as the mean over all Structural connectomes).
In Figure 4b, we show the equi-M set for the prototypical Structural connectome, computed as the mean of all Structural connectomes in the data set (group average). This is a different way of checking that, while details in the size and shape of equi-M sets vary across connectomes, the presence of the cusp on the right appears to be a robust feature for Structural connectomes, for all individual and group representations.
@@
As described in the previous paragraphs, we calculated mutual Pearson correlations between the measures in each category, but also across categories. This allows us to evaluate to what extent the architecture of the Structural connectome can predict patterns in asymptotic dynamics, as reflected in the geometric properties of the equi-M set. Figure 5 describes all these correlations in a table in which the colors represent positive and negative correlation values.
@@
Figure 5: Correlations between the graph-theoretical measures and geometric measures for the Structural connectomes (left) and for randomizations of the Structural connectomes that preserve the number of edges into each node (right). The measures are listed along both coordinate axes, in the order found in the tables, and in the order introduced in the text. For $C$, $\operatorname{Re}(R)$ and $\operatorname{Re}(L)$, we used their absolute values, to illustrate the distance to the origin. For the complex modes $m_{-2}$ to $m_{2}$, we use their moduli (to illustrate their strength). In this scheme, the three diagonal squares show correlations between measures within the same category, and the off-diagonal rectangles show correlations between measures in different categories. For simplicity, we only included the $\rho$ values here.
@@
To reiterate, let us notice that there is a significantly positively-correlated block of graph-theoretical measures (all except $BC$ and $A$) – we call this block $\mathcal{G}$. There is also a highly correlated block of geometric measures (all except $\epsilon$) – we call this block $\mathcal{T}$. Finally, we noted that there is also some degree of correlation within the collection of seven Fourier-derived measures (less significant as modes increase). In addition to this, Figure 5 shows that there are strong negative correlations between the measures in $\mathcal{G}$ and the measures in $\mathcal{T}$. The only ones that fail the significance level $\alpha=0.01$ are the correlations with the average Local Efficiency $\langle LE\rangle$. Since the geometric measures in $\mathcal{G}$ all relate in some way to the size of the set, this result supports an idea that we had previously discussed in our work on both theoretical and Structural connectomes: higher network wiring connectivity and efficiency (higher node degree and centrality, stronger motifs) are associated with smaller equi-M sets. In turn, the graph measures in $\mathcal{G}$ are also negatively correlated with the Fourier measures, with the most significant values for the centroid distance $|G|$ and the perimeter $P$ (which is not unexpected, since these are also measures of “size” and show strong positive correlations with the other measures of size in $\mathcal{T}$). What is interesting is that the moduli of the modes also show similar correlation patterns, even though the strength of the correlations weakens with increasing mode index (and significance decreases). This suggests a tie between the wiring of the underlying graph and the geometry and variability of the boundary of the equi-M set, tie which transcends simple size-related properties for Structural connectomes. We will revisit this point for perspective, after we discuss the same relationships in Functional connectomes in Section 3.2.
@@
Table 3: Network measure statistics for randomized Structural connectomes: Betweenness Centrality (BC); Assortativity (AS); Clustering Coefficient (CC); Degree (D); Global efficiency (GE); Local efficiency (LE); Eigencentrality (EC); Transitivity (T); Mean 3D Motifs 1 (M1) and 2 (M2). For the node-wise measures, the mean was first computed over all nodes. Then, means and standard deviations over the whole subject population are shown in the table for each of the measures used. The strengths of the correlations between these measures ($\rho$ values for Pearson correlations) are illustrated in Figure 5.
@@
Mode $m_{-2}$ $12.20-4.64i$ 10.27
Mode $m_{-1}$ $-128.93+21,60i$ 34.43
Mode $m_{0}$ $-67.18+0.09i$ 21.42
Mode $m_{1}$ $-14.92-3.14i$ 9.58
Mode $m_{2}$ $-19.39-6.98i$ 9.27
@@
Finally, let us point out that Betweenness Centrality and Assortativity don’t have any significant correlations with any of the equi-M measures we used. The natural next question is whether any one of the strongly correlated measures in $\mathcal{G}$ holds the key to the shape of the equi-M set. To explore this question further, we simulated an in silico experiment aimed to investigate to what extent the average node Degree dictates in and off itself the measures of the equi-M set. To do so, we chose a random Structural connectome (subject #2), and generated a set of 47 additional surrogate connectomes with the same node degree distribution as the original. The 48 resulting sets are shown in Figure 6 (as contours), together with the prototypical set for the group of random connectomes. The statistics of geometric and Fourier measures are shown in Tables 3 and 4 LABEL:structural_random_graph_measures,structural_random_geometric_measures.
Figure 6: Visualizations of equi-M set statistics for the randomly generated Structural connectome data. (a) Statistical equi-M set (the light to dark gradient shows the fraction of equi-M sets that the respective point in $\mathbb{C}$ belongs to), together with the the collection of contours of all equi-M sets derived from the randomly generated Structural connectomes (overlaid in gray). (b) The equi-M set computed for the corresponding prototypical connectome.
Notice that the random networks created this way are quite different from the original ones, with different graph-theoretical statistics. They exhibit collectively narrower spread, more generally than just for the Degree, and with distinct global properties (e.g., they are slightly disassortative). In turn, the landmark landscape of the corresponding equi-M sets is different, and the labels generally identify larger sets (as visible from the contours in Figure 6a), and with significantly more information contained in the first two modes. Interestingly, notice that the eccentricities are comparable to those in the original distribution, suggesting that eccentricity may be robust to Degree-preserving alterations of the network. Another interesting fact to note is that the prototypical equi-M set in this case is a rescaled version of the traditional Mandelbrot set. This may reflect eliminating the randomness of the network by averaging over connectomes, with the resulting system behaving equivalently to a single node via a multiplicative weight.
Finally, notice that, for randomly generated connectomes (unlike for the data-driven Structural connectomes) there are no significant correlations between the graph measures and the geometric/Fourier measures (as seen in Figure 5b). This suggests that the tie between the structure and the network’s asymptotic dynamics may not simply be intrinsic, or achieved by chance, but rather the reflection of the natural design in functional systems.
@@ 3.2 Functional connectomes
From Section 2.1, we consider two types of Functional connectomes: one corresponding to a scan with no task (Rest) and one for an Emotional task. We first describe the graph-theoretical characteristics and the geometry of the equi-M sets for both tasks separately. Then, we will compute correlations between them and draw comparisons between the results, as well as with those obtained for the Structural connectomes. As discussed in Section 2.2, it is common practice in network science to describe signed connectomes by separating their positive and negative sub-networks. Then, we can produce graph-theoretical descriptions for each sub-network, which are used as a characterization of the overall, signed network. In order to carry on our study of the relationship between network architecture and dynamics, we compute graph-theoretical measures for the positive sub-network, the negative sub-networks and the absolute value of each connectome. Our analysis of Structural connectomes concluded that, in the case of positive symmetric connectomes, graph-theoretical measures are predictive of geometric measures. A crucial part of our investigation of signed connectomes is to establish whether these partial representations of the network, taken together, can perform the same predictive function or if the information they provide is insufficient. If it is not sufficient, we can conclude that a crucial part of the emerging dynamics relies on the way in which they interact that is not captured in these partial measures.
@@ 3.2.1 Measures for Functional connectomes
For Rest and Emotion connectomes, graph theoretical measures are described in Tables 5 and 6. Further statistical differences between these sets of measures are described and interpreted in Section 3.2.3.
Table 5: Network measure statistics for Rest connectomes. Graph-theoretical measures were computed separately for the positive and for the negative sub-networks of the connectome, as well as for the absolute value of the connectome: Betweenness Centrality (BC); Assortativity (AS); Clustering Coefficient (CC); Degree (D); Global efficiency (GE); Local efficiency (LE); Eigencentrality (EC); Transitivity (T); Mean 3D Motifs 1 (M1) and 2 (M2). For the node-wise measures, the mean was first computed over all nodes. Then mean $\mu$ and standard deviation $\sigma$ over the whole subject population of each of the measures used are shown in the table.
@@
$BC$ $303.22$ $11.21$ $174.11$ $8.42$ $157.46$ $9.44$
$AS$ $0.06$ $0.03$ $-0.06$ $0.02$ $-0.05$ $<0.001$
$\langle EC\rangle$ $0.06$ $0.004$ $0.06$ $0.008$ $0.06$ $<0.001$
@@
Table 6: Network measure statistics for Emotion connectomes. Graph-theoretical measures were computed separately for the positive and for the negative sub-networks of the connectome, as well as for the absolute value of the connectome. Betweenness Centrality (BC); Assortativity (AS); Clustering Coefficient (CC); Degree (D); Global efficiency (GE); Local efficiency (LE); Eigencentrality (EC); Transitivity (T); Mean 3D Motifs 1 (M1) and 2 (M2). For the node-wise measures, the mean was first computed over all nodes. Then mean $\mu$ and standard deviation $\sigma$ over the whole subject population of each of the measures used are shown in the table.
@@
Geometric measures were calculated for all Rest connectomes and are shown in Tables 7 and 8.
Table 7: Geometric statistics for Functional connectomes. The mean $\mu$ and standard deviation $\sigma$ over the whole subject population were computed for each one-dimensional measure, for the Rest connectomes (top) and the Emotion connectomes (bottom). The strengths of the correlations between these measures ($\rho$ values for Pearson correlations) are illustrated in Figure 10.
@@
Table 8: Fourier-based measure statistics for Functional connectomes. The mean $\mu$ and standard deviation $\sigma$ over the whole subject population were computed for each one-dimensional measure. For the modes (which are distributions in $\mathbb{C}$), more information on standard deviations and principal components is included later in the section. The strengths of the correlations between these measures ($\rho$ values for Pearson correlations) are illustrated in Figure 10.

### v2

@@ 3.1 Structural connectomes
The graph-theoretical measures described in Section˜2.2 were computed for all Structural connectomes. The mean and standard deviation for each of these measures are provided in Table˜1. In Figure˜5, we have the pairwise correlation between the graph-theoretical measures. Notice that, aside from $BC$ and $AS$, the other network measures show significant pairwise positive correlations with significance values $p<0.0007$.
Table 1: Network measures statistics for Structural connectomes: Betweenness Centrality (BC), Assortativity (AS), Clustering Coefficient ($C$), Degree (D), Global Efficiency (GE), Local Efficiency (LE), Eigencentrality (EC), Transitivity (TR), Mean 3D Motifs 1 (M1) and 2 (M2). For the node-wise measures, the mean was first computed over all nodes. Then, mean $\mu$ and standard deviation $\sigma$ over the whole subject population are shown in the table for each of the measures used. The strengths of the correlations between these measures ($\rho$ values for Pearson correlations) are illustrated in Figure˜5.
@@
The measures described in Section˜2.4 were calculated for each Structural connectome. Their mean and standard deviation are reported in Table˜2. To start with, Figure˜12 readily conveys certain unifying and differentiating features across the Structural equi-M sets. First, we notice relatively broad differences in the size of the sets between subjects. These differences reflect relatively broad distributions in the values of the geometric landmarks and distances (for all diameters $d_{XY}$ that we calculated, the standard deviation is at least half of the mean). The one measure which has a significantly smaller relative variance is the ratio $\epsilon$. This is because the “eccentricity” of the sets remains relatively robust across the connectomes, with eccentricities tightly clustered around the average $\mu(\epsilon)=0.81$, regardless of size.
Table 2: Geometric measure statistics for Structural connectomes. Then mean $\mu$ and standard deviation $\sigma$ over the whole subject population were computed for each geometric measure. For the modes, the standard deviation $\sigma$ was computed by treating the real and imaginary parts as independent and calculating the standard deviation as the square root of the sum of the variances of the real and imaginary parts. The strengths of the correlations between these measures ($\rho$ values for Pearson correlations) are illustrated in Figure˜5.
@@
Mode $m_{-2}$ $7.79$ 2.59-2.59 $6.33$
Mode $m_{-1}$ 66.97-66.97 $10.56$ $32.28$
Mode $m_{0}$ 39.05-39.05 0.09-0.09 $20.15$
Mode $m_{1}$ 10.52-10.52 1.79-1.79 $6.38$
Mode $m_{2}$ 9.42-9.42 3.15-3.15 $6.19$
@@
As discussed in Section˜2.3, we approach the collection of sets obtained from connectome data in a statistical way by of considering the distribution of equi-M sets over all connectomes under consideration. We did this by computing the equi-M set for each individual connectome and assigning to each point in the complex plane the fraction of individual equi-M sets that contain the point (effectively measuring how “likely” it is for the given point to be in any equi-M set generated by our data). Figure˜4a illustrates this statistical equi-M set with the fraction represented by the color intensity, increasing from light to dark. To emphasize the underpinnings of this representation, we overlaid on it the contours of all equi-M sets for Structural connectomes (in gray), outlining the basis of how these statistics were computed. Consistent with our previous comment, all these contours exhibit a cusp on the right side along the real axis. Note, in particular, that the black center region is the $c$-locus that is common to the equi-M sets for all Structural connectomes.
Figure 4: Different approaches to the Structural connectome data using the equi-M set. (a) Statistical equi-M set: the light to dark gradient shows the fraction of equi-M sets that the respective point in $\mathbb{C}$ belongs to. (b) The equi-M set computed for the prototypical Structural connectome (i.e., the matrix computed as the mean over all Structural connectomes).
In Figure˜4b, we show the equi-M set for the prototypical Structural connectome, computed as the mean of all Structural connectomes in the data set (group average). This is a different way of checking that, while details in the size and shape of equi-M sets vary across connectomes, the presence of the cusp on the right appears to be a robust feature for Structural connectomes, for all individual and group representations.
@@
As described in the previous paragraphs, we calculated mutual Pearson correlations between the measures in each category, but also across categories. This allows us to evaluate to what extent the architecture of the Structural connectome can predict patterns in asymptotic dynamics, as reflected in the geometric properties of the equi-M set. Figure˜5 describes all these correlations in a table in which the colors represent positive and negative correlation values.
@@
Figure 5: Correlations between the graph-theoretical measures and geometric measures for the Structural connectomes (panel A) and for randomizations of the Structural connectomes that preserve the number of edges into each node (panel B) . The measures are listed along both coordinate axes, in the order found in the tables, and in the order introduced in the text. For $C$, $\operatorname{Re}(R)$ and $\operatorname{Re}(L)$, we used their absolute values, to illustrate the distance to the origin. For the complex modes $m_{-2}$ to $m_{2}$, we use their moduli (to illustrate their strength). In this scheme, the three diagonal squares show correlations between measures within the same category, and the off-diagonal rectangles show correlations between measures in different categories. For simplicity, we only included the $\rho$ values here.
@@
To reiterate, let us notice that there is a significantly positively-correlated block of graph-theoretical measures (all except $BC$ and $AS$) – we call this block $\mathcal{G}$. There is also a highly correlated block of geometric measures (all except $\epsilon$) – we call this block $\mathcal{T}$. Finally, we noted that there is also some degree of correlation within the collection of seven Fourier-derived measures (less significant as modes increase). In addition to this, Figure˜5 shows that there are strong negative correlations between the measures in $\mathcal{G}$ and the measures in $\mathcal{T}$. The only ones that fail the significance level $\alpha=0.01$ are the correlations with the average Local Efficiency $\langle LE\rangle$. Since the geometric measures in $\mathcal{G}$ all relate in some way to the size of the set, this result supports an idea that we had previously discussed in our work on both theoretical and Structural connectomes: higher network wiring connectivity and efficiency (higher node degree and centrality, stronger motifs) are associated with smaller equi-M sets. In turn, the graph measures in $\mathcal{G}$ are also negatively correlated with the Fourier measures, with the most significant values for the centroid distance $|G|$ and the perimeter $P$ (which is not unexpected, since these are also measures of “size” and show strong positive correlations with the other measures of size in $\mathcal{T}$). What is interesting is that the moduli of the modes also show similar correlation patterns, even though the strength of the correlations weakens with increasing mode index (and significance decreases). This suggests a tie between the wiring of the underlying graph and the geometry and variability of the boundary of the equi-M set, tie which transcends simple size-related properties for Structural connectomes. We will revisit this point for perspective, after we discuss the same relationships in Functional connectomes in Section˜3.2.
@@
Table 3: Network measure statistics for randomized Structural connectomes: Betweenness Centrality (BC); Assortativity (AS); Clustering Coefficient ($C$); Degree (D); Global efficiency (GE); Local efficiency (LE); Eigencentrality (EC); Transitivity (TR); Mean 3D Motifs 1 (M1) and 2 (M2). For the node-wise measures, the mean was first computed over all nodes. Then, means and standard deviations over the whole subject population are shown in the table for each of the measures used. The strengths of the correlations between these measures ($\rho$ values for Pearson correlations) are illustrated in Figure 5.
@@
Mode $m_{-2}$ 12.204.64i12.20-4.64i 10.27
Mode $m_{-1}$ 128.93+21,60i-128.93+21,60i 34.43
Mode $m_{0}$ 67.18+0.09i-67.18+0.09i 21.42
Mode $m_{1}$ 14.923.14i-14.92-3.14i 9.58
Mode $m_{2}$ 19.396.98i-19.39-6.98i 9.27
@@
Finally, let us point out that Betweenness Centrality and Assortativity don’t have any significant correlations with any of the equi-M measures we used. The natural next question is whether any one of the strongly correlated measures in $\mathcal{G}$ holds the key to the shape of the equi-M set. To explore this question further, we simulated an in silico experiment aimed to investigate to what extent the average node Degree dictates in and off itself the measures of the equi-M set. To do so, we chose a random Structural connectome (subject #2), and generated a set of 47 additional surrogate connectomes with the same node degree distribution as the original. The 48 resulting sets are shown in Figure˜6 (as contours), together with the prototypical set for the group of random connectomes. The statistics of geometric and Fourier measures are shown in Tables 3 and 4.
Figure 6: Visualizations of equi-M set statistics for the randomly generated Structural connectome data. (a) Statistical equi-M set: the light to dark gradient shows the fraction of equi-M sets that the respective point in $\mathbb{C}$ belongs to. (b) The equi-M set computed for the corresponding prototypical connectome.
Notice that the random networks created this way are quite different from the original ones, with different graph-theoretical statistics. They exhibit collectively narrower spread, more generally than just for the Degree, and with distinct global properties (e.g., they are slightly disassortative). In turn, the landmark landscape of the corresponding equi-M sets is different, and the labels generally identify larger sets (as visible from the contours in Figure˜6a), and with significantly more information contained in the first two modes. Interestingly, notice that the eccentricities are comparable to those in the original distribution, suggesting that eccentricity may be robust to Degree-preserving alterations of the network. Another interesting fact to note is that the prototypical equi-M set in this case is a rescaled version of the traditional Mandelbrot set. This may reflect eliminating the randomness of the network by averaging over connectomes, with the resulting system behaving equivalently to a single node via a multiplicative weight.
Finally, notice that, for randomly generated connectomes (unlike for the data-driven Structural connectomes) there are no significant correlations between the graph measures and the geometric/Fourier measures (as seen in Figure˜5b). This suggests that the tie between the structure and the network’s asymptotic dynamics may not simply be intrinsic, or achieved by chance, but rather the reflection of the natural design in functional systems.
@@ 3.2 Functional connectomes
From Section˜2.1, we consider two types of Functional connectomes: one corresponding to a scan with no task (Rest) and one for an Emotional task. We first describe the graph-theoretical characteristics and the geometry of the equi-M sets for both tasks separately. Then, we will compute correlations between them and draw comparisons between the results, as well as with those obtained for the Structural connectomes. As discussed in Section˜2.2, it is common practice in network science to describe signed connectomes by separating their positive and negative sub-networks. Then, we can produce graph-theoretical descriptions for each sub-network, which are used as a characterization of the overall, signed network. In order to carry on our study of the relationship between network architecture and dynamics, we compute graph-theoretical measures for the positive sub-network, the negative sub-networks and the absolute value of each connectome. Our analysis of Structural connectomes concluded that, in the case of positive symmetric connectomes, graph-theoretical measures are predictive of geometric measures. A crucial part of our investigation of signed connectomes is to establish whether these partial representations of the network, taken together, can perform the same predictive function or if the information they provide is insufficient. If it is not sufficient, we can conclude that a crucial part of the emerging dynamics relies on the way in which they interact that is not captured in these partial measures.
@@ 3.2.1 Measures for Functional connectomes
For Rest and Emotion connectomes, graph theoretical measures are described in Tables˜5 and 6. As discussed in the introduction, these measures are not available for signed networks. We therefore computed them for the associated networks, as done in a typical network study. The table presents the mean and standard deviations for each measure, for each associated network. Further statistical differences between these sets of measures are described and interpreted in Section˜3.2.3.
Table 5: Network measure statistics for Rest connectomes. Graph-theoretical measures were computed separately for the positive and for the negative sub-networks of the connectome, as well as for the absolute value of the connectome: Betweenness Centrality (BC); Assortativity (AS); Clustering Coefficient ($C$); Degree (D); Global efficiency (GE); Local efficiency (LE); Eigencentrality (EC); Transitivity (TR); Mean 3D Motifs 1 (M1) and 2 (M2). For the node-wise measures, the mean was first computed over all nodes. Then mean $\mu$ and standard deviation $\sigma$ over the whole subject population of each of the measures used are shown in the table.
@@
BCBC $303.22$ $11.21$ $174.11$ $8.42$ $157.46$ $9.44$
ASAS $0.06$ $0.03$ 0.06-0.06 $0.02$ 0.05-0.05 <0.001<0.001
$\langle EC\rangle$ $0.06$ $0.004$ $0.06$ $0.008$ $0.06$ <0.001<0.001
@@
Table 6: Network measure statistics for Emotion connectomes. Graph-theoretical measures were computed separately for the positive and for the negative sub-networks of the connectome, as well as for the absolute value of the connectome. Betweenness Centrality (BC); Assortativity (AS); Clustering Coefficient ($C$); Degree (D); Global efficiency (GE); Local efficiency (LE); Eigencentrality (EC); Transitivity (TR); Mean 3D Motifs 1 (M1) and 2 (M2). For the node-wise measures, the mean was first computed over all nodes. Then mean $\mu$ and standard deviation $\sigma$ over the whole subject population of each of the measures used are shown in the table.
@@
Geometric measures were calculated for all Rest and Emotion connectomes and are shown in Tables˜7 and 8. In the table, we show only measures for the full, signed network, since one of our goals is precisely to establish to what extent partial graph theoretical information from the associate networks is relevant to the whole-network dynamics. To support this choice, we also computed the equi-M sets for the associated networks for all Functional connectomes (Rest and Emotion). What we observed is that these sets were virtually all scaled versions of the traditional Mandelbrot set (see for example Figures 14 and 16 in the Appendix for the equi-M sets corresponding to the absolute value connectomes). This suggests that our intuition is correct, and that the associate networks cannot recreate, taken separately, the complexity of the overall network dynamics. This will be further discussed in the following sections.
Table 7: Geometric statistics for Functional connectomes. The mean $\mu$ and standard deviation $\sigma$ over the whole subject population were computed for each one-dimensional measure, for the Rest connectomes (top) and the Emotion connectomes (bottom). The strengths of the correlations between these measures ($\rho$ values for Pearson correlations) are illustrated in Figure˜10.
@@
Table 8: Fourier-based measure statistics for Functional connectomes. The mean $\mu$ and standard deviation $\sigma$ over the whole subject population were computed for each one-dimensional measure. For the modes (which are distributions in $\mathbb{C}$), more information on standard deviations and principal components is included later in the section. The strengths of the correlations between these measures ($\rho$ values for Pearson correlations) are illustrated in Figure˜10.